OpenStudy (anonymous):

marc is growing three plants plant A is 1/3 plant B is 3/4 plant C is 5/6 order the plants from smallest to tallest <3

OpenStudy (anonymous):

ok so multiply all of the denominators to a greatest common factor. It would be 12 so 1/3(4) = 4/12 3/4(3) 9/12 5/6(2)= 10/12 so plant c,b,a get it?

OpenStudy (anonymous):

oh wait a , b ,c thats smallest to largest lol

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

best response itttt (: